# Origin of effective graviton mass from spontaneous symmetry breaking of scalar field

### v1

@@ Appendix B Effective graviton mass generation in arbitrary curved spacetime
where $\bar{g}_{\mu\nu}$ is the metric describing the background geometry w.r.t which we linearised the theory of gravity, $h_{\mu\nu}$ is the metric perturbation and $\varepsilon$ carries the order of the metric perturbation. This leads to the following expression of the inverse metric:
@@
$\begin{split}\Gamma_{\ \mu\nu}^{\rho}&=\bar{\Gamma}_{\ \mu\nu}^{\rho}+\varepsilon\Big{[}\frac{1}{2}\bar{g}^{\rho\lambda}(\partial_{\mu}h_{\lambda\nu}+\partial_{\nu}h_{\mu\lambda}-\partial_{\lambda}h_{\mu\nu})-\bar{g}^{\rho\alpha}h_{\alpha\beta}\bar{\Gamma}_{\ \mu\nu}^{\beta}\Big{]}\\ &-\varepsilon^{2}\Big{[}\frac{1}{2}\bar{g}^{\beta\lambda}h_{\ \beta}^{\rho}(\partial_{\mu}h_{\lambda\nu}+\partial_{\nu}h_{\mu\lambda}-\partial_{\lambda}h_{\mu\nu})-\bar{g}^{\alpha\beta}h_{\ \beta}^{\rho}h_{\alpha\sigma}\bar{\Gamma}_{\ \mu\nu}^{\sigma}\Big{]},\end{split}$ (79)
@@
$\begin{split}*&=(\Gamma_{\ \sigma\nu}^{\beta})_{L}\Big{[}\bar{\nabla}_{\rho}(\bar{g}^{\mu\alpha}h_{\alpha\beta})-\frac{1}{2}\bar{g}^{\mu\alpha}(\bar{\nabla}_{\rho}h_{\alpha\beta}+\bar{\nabla}_{\beta}h_{\rho\alpha}-\bar{\nabla}_{\alpha}h_{\rho\beta})\Big{]}\\ &=(\Gamma_{\ \sigma\nu}^{\beta})_{L}\Big{[}\frac{1}{2}\bar{g}^{\mu\alpha}(\bar{\nabla}_{\rho}h_{\alpha\beta}+\bar{\nabla}_{\alpha}h_{\rho\beta}-\bar{\nabla}_{\beta}h_{\rho\alpha})\Big{]}\\ &=(\Gamma_{\ \sigma\nu}^{\beta})_{L}\Big{[}\frac{1}{2}\bar{g}^{\mu\alpha}\delta_{\beta}^{\ \lambda}(\bar{\nabla}_{\rho}h_{\alpha\lambda}+\bar{\nabla}_{\alpha}h_{\rho\lambda}-\bar{\nabla}_{\lambda}h_{\rho\alpha})\Big{]}\\ &=(\Gamma_{\ \sigma\nu}^{\beta})_{L}\Big{[}\frac{1}{2}\bar{g}^{\mu\alpha}\bar{g}_{\beta\gamma}\bar{g}^{\gamma\lambda}(\bar{\nabla}_{\rho}h_{\alpha\lambda}+\bar{\nabla}_{\alpha}h_{\rho\lambda}-\bar{\nabla}_{\lambda}h_{\rho\alpha})\Big{]}\\ &=\bar{g}^{\mu\alpha}\bar{g}_{\beta\gamma}(\Gamma_{\ \sigma\nu}^{\beta})_{L}(\Gamma_{\ \rho\alpha}^{\gamma})_{L}.\end{split}$ (97)
@@
$\begin{split}R&=\bar{R}+\varepsilon[\bar{g}^{\nu\sigma}(R_{\nu\sigma})_{L}-h^{\nu\sigma}\bar{R}_{\nu\sigma}]-\varepsilon^{2}\{\bar{g}^{\nu\sigma}h_{\ \beta}^{\mu}(R_{\ \nu\mu\sigma}^{\beta})_{L}+h^{\nu\sigma}(R_{\nu\sigma})_{L}\\ &-h^{\nu\lambda}h_{\ \lambda}^{\sigma}\bar{R}_{\nu\sigma}+\bar{g}^{\nu\sigma}\bar{g}^{\mu\alpha}\bar{g}_{\beta\gamma}[(\Gamma_{\ \sigma\nu}^{\beta})_{L}(\Gamma_{\ \mu\alpha}^{\gamma})_{L}-(\Gamma_{\ \mu\nu}^{\beta})_{L}(\Gamma_{\ \sigma\alpha}^{\gamma})_{L}]\}\\ &=\bar{R}+\varepsilon\Big{[}\frac{1}{2}\bar{g}^{\nu\sigma}(\bar{\nabla}_{\mu}\bar{\nabla}_{\sigma}h_{\ \nu}^{\mu}+\bar{\nabla}_{\mu}\bar{\nabla}_{\nu}h_{\ \sigma}^{\mu}-\bar{\Box}h_{\sigma\nu}-\bar{\nabla}_{\sigma}\bar{\nabla}_{\nu}h)-h^{\nu\sigma}\bar{R}_{\nu\sigma}\Big{]}\\ &-\varepsilon^{2}\Bigg{[}\frac{1}{2}\bar{g}^{\nu\sigma}h_{\ \mu}^{\rho}[\bar{\nabla}_{\rho}\bar{\nabla}_{\sigma}h_{\ \nu}^{\mu}+\bar{\nabla}_{\rho}\bar{\nabla}_{\nu}h_{\ \sigma}^{\mu}-\bar{\nabla}_{\rho}\bar{\nabla}^{\mu}h_{\sigma\nu}-\bar{\nabla}_{\sigma}\bar{\nabla}_{\rho}h_{\ \nu}^{\mu}-\bar{\nabla}_{\sigma}\bar{\nabla}_{\nu}h_{\ \rho}^{\mu}+\bar{\nabla}_{\sigma}\bar{\nabla}^{\mu}h_{\rho\nu}]\\ &+\frac{1}{2}h^{\nu\sigma}(\bar{\nabla}_{\mu}\bar{\nabla}_{\sigma}h_{\ \nu}^{\mu}+\bar{\nabla}_{\mu}\bar{\nabla}_{\nu}h_{\ \sigma}^{\mu}-\bar{\Box}h_{\sigma\nu}-\bar{\nabla}_{\sigma}\bar{\nabla}_{\nu}h)\\ &-h^{\nu\lambda}h_{\ \lambda}^{\sigma}\bar{R}_{\nu\sigma}+\frac{1}{4}\Big{[}4\bar{\nabla}_{\sigma}h_{\rho}^{\ \sigma}\bar{\nabla}_{\mu}h^{\rho\mu}+2\bar{\nabla}^{\mu}h^{\lambda\sigma}\bar{\nabla}_{\lambda}h_{\sigma\mu}-4\bar{\nabla}_{\sigma}h^{\lambda\sigma}\bar{\nabla}_{\lambda}h+\bar{\nabla}_{\rho}h\bar{\nabla}^{\rho}h-3\bar{\nabla}_{\rho}h_{\mu\nu}\bar{\nabla}^{\rho}h^{\mu\nu}\Big{]}\Bigg{]}.\end{split}$ (106)
@@
$\begin{split}\sqrt{-g}R&=\sqrt{-\bar{g}}\Bigg{[}\bar{R}+\varepsilon\underbrace{\left(\frac{1}{2}h\bar{R}-h^{\nu\sigma}\bar{R}_{\nu\sigma}\right)}_{=0}+\varepsilon\underbrace{\Big{[}\bar{\nabla}_{\mu}\bar{\nabla}_{\nu}h^{\mu\nu}-\bar{\Box}h\Big{]}}_{\text{surface terms}}+\frac{\varepsilon^{2}}{2}h\Big{[}\bar{\nabla}_{\mu}\bar{\nabla}_{\nu}h^{\mu\nu}-\bar{\Box}h\Big{]}\\ &-\frac{\varepsilon^{2}}{4}\left(h_{\alpha\beta}h^{\alpha\beta}-\frac{1}{2}h^{2}\right)\bar{R}+\varepsilon^{2}\bar{R}_{\nu\sigma}\left(h^{\nu\lambda}h_{\ \lambda}^{\sigma}-\frac{1}{2}hh^{\nu\sigma}\right)\\ &-\varepsilon^{2}\left(\frac{1}{4}\bar{\nabla}_{\lambda}h_{\mu\nu}\bar{\nabla}^{\lambda}h^{\mu\nu}+\frac{1}{4}\bar{\nabla}_{\rho}h\bar{\nabla}^{\rho}h-\frac{1}{2}\bar{\nabla}^{\nu}h^{\rho\mu}\bar{\nabla}_{\mu}h_{\rho\nu}\right)\Bigg{]}\\ &=\sqrt{-\bar{g}}\Bigg{[}\bar{R}-\frac{\varepsilon^{2}}{4}\left(h_{\alpha\beta}h^{\alpha\beta}-\frac{1}{2}h^{2}\right)\bar{R}+\varepsilon^{2}\bar{R}_{\nu\sigma}\left(h^{\nu\lambda}h_{\ \lambda}^{\sigma}-\frac{1}{2}hh^{\nu\sigma}\right)\\ &-\varepsilon^{2}\left(\frac{1}{4}\bar{\nabla}_{\lambda}h_{\mu\nu}\bar{\nabla}^{\lambda}h^{\mu\nu}-\frac{1}{4}\bar{\nabla}_{\rho}h\bar{\nabla}^{\rho}h-\frac{1}{2}\bar{\nabla}^{\nu}h^{\rho\mu}\bar{\nabla}_{\mu}h_{\rho\nu}+\frac{1}{2}\bar{\nabla}_{\mu}h\bar{\nabla}_{\nu}h^{\mu\nu}\right)\Bigg{]}.\end{split}$ (108)
Now considering $\varepsilon=\kappa$ and rescaling the action of gravitons by a factor of 4 (this is equivalent to taking $\varepsilon=2\kappa$), the linearised Einstein-Hilbert action in arbitrary curved spacetime can be expressed as
$\begin{split}S_{EH}&=\frac{1}{2\kappa^{2}}\int d^{4}x\sqrt{-g}R=\frac{1}{2\kappa^{2}}\int d^{4}x\sqrt{-\bar{g}}\bar{R}\\ &+\int d^{4}x\sqrt{-\bar{g}}\Big{[}-\frac{1}{2}\bar{\nabla}_{\lambda}h_{\mu\nu}\bar{\nabla}^{\lambda}h^{\mu\nu}+\frac{1}{2}\bar{\nabla}_{\rho}h\bar{\nabla}^{\rho}h+\bar{\nabla}^{\nu}h^{\rho\mu}\bar{\nabla}_{\mu}h_{\rho\nu}-\bar{\nabla}_{\mu}h\bar{\nabla}_{\nu}h^{\mu\nu}\\ &-\frac{1}{2}\bar{R}\left(h_{\alpha\beta}h^{\alpha\beta}-\frac{1}{2}h^{2}\right)+2\bar{R}_{\nu\sigma}\left(h^{\nu\lambda}h_{\ \lambda}^{\sigma}-\frac{1}{2}hh^{\nu\sigma}\right)\Big{]}.\end{split}$ (109)
@@
$\begin{split}S_{M}[\Phi,\Phi^{*},g_{\mu\nu}]&=S_{M}[\Phi,\Phi^{*},\bar{g}_{\mu\nu}]+\frac{\kappa}{2}\int\sqrt{-\bar{g}}h_{\mu\nu}T^{\mu\nu}-\frac{\kappa^{2}}{4}\int d^{4}x\sqrt{-\bar{g}}h^{\mu\nu}h_{\rho\sigma}\mathscr{O}_{\mu\nu}^{\ \ \rho\sigma}+\ldots.\end{split}$ (110)
We may note that the linear part cancels with the second term in the first line of the equation (108) since this is nothing but the Einstein’s equation in the presence of matter source. Moreover, it is mathematically consistent since if we expand the total matter + E-H action around a solution of classical equation of motion, the linear terms drop out and leading order terms are quadratic order in the perturbations. If we rescale again the graviton contribution by a factor of $4$ coming from the matter sector, we obtain the following total action in the gravitons in curved spacetime
$\begin{split}S_{h}&=\int d^{4}x\sqrt{-\bar{g}}\Big{[}-\frac{1}{2}\bar{\nabla}_{\lambda}h_{\mu\nu}\bar{\nabla}^{\lambda}h^{\mu\nu}+\frac{1}{2}\bar{\nabla}_{\rho}h\bar{\nabla}^{\rho}h+\bar{\nabla}^{\nu}h^{\rho\mu}\bar{\nabla}_{\mu}h_{\rho\nu}-\bar{\nabla}_{\mu}h\bar{\nabla}_{\nu}h^{\mu\nu}\\ &-\frac{1}{2}\bar{R}\left(h_{\alpha\beta}h^{\alpha\beta}-\frac{1}{2}h^{2}\right)+2\bar{R}_{\nu\sigma}\left(h^{\nu\lambda}h_{\ \lambda}^{\sigma}-\frac{1}{2}hh^{\nu\sigma}\right)\Big{]}-\kappa^{2}\int d^{4}x\sqrt{-\bar{g}}h^{\mu\nu}h_{\rho\sigma}\mathscr{O}_{\mu\nu}^{\ \ \rho\sigma},\end{split}$ (111)
@@
where $V(|\Phi_{0}|)$ turns out to be negative after spontaneous symmetry breaking of the field potential. Substituting all the above information including the fact $\bar{R}=-\kappa^{2}T$ and $\bar{R}_{\nu\sigma}=\kappa^{2}\left(T_{\nu\sigma}-\frac{1}{2}\bar{g}_{\nu\sigma}T\right)$ where $T=-4V(|\Phi_{0}|)$ and $T_{\nu\sigma}=-\bar{g}_{\nu\sigma}V(|\Phi_{0}|)$ in the leading order, we obtain the following quadratic action of gravitons
$\begin{split}S_{h}&=\int d^{4}x\sqrt{-\bar{g}}\Big{[}-\frac{1}{2}\bar{\nabla}_{\lambda}h_{\mu\nu}\bar{\nabla}^{\lambda}h^{\mu\nu}+\frac{1}{2}\bar{\nabla}_{\rho}h\bar{\nabla}^{\rho}h+\bar{\nabla}^{\nu}h^{\rho\mu}\bar{\nabla}_{\mu}h_{\rho\nu}-\bar{\nabla}_{\mu}h\bar{\nabla}_{\nu}h^{\mu\nu}\\ &-2\kappa^{2}V(|\Phi_{0}|)\left(h_{\alpha\beta}h^{\alpha\beta}-\frac{1}{2}h^{2}\right)+2\kappa^{2}V(|\Phi_{0}|)\left(h_{\alpha\beta}h^{\alpha\beta}-\frac{1}{2}h^{2}\right)\Big{]}\\ &+\kappa^{2}V(|\Phi_{0}|)\int d^{4}x\sqrt{-\bar{g}}\left(h_{\mu\nu}h^{\mu\nu}-\frac{1}{2}h^{2}\right)\\ &=\int d^{4}x\sqrt{-\bar{g}}\Big{[}-\frac{1}{2}\bar{\nabla}_{\lambda}h_{\mu\nu}\bar{\nabla}^{\lambda}h^{\mu\nu}+\frac{1}{2}\bar{\nabla}_{\rho}h\bar{\nabla}^{\rho}h+\bar{\nabla}^{\nu}h^{\rho\mu}\bar{\nabla}_{\mu}h_{\rho\nu}-\bar{\nabla}_{\mu}h\bar{\nabla}_{\nu}h^{\mu\nu}\\ &+\kappa^{2}V(|\Phi_{0}|)\left(h_{\alpha\beta}h^{\alpha\beta}-\frac{1}{2}h^{2}\right)\Big{]}.\end{split}$ (116)
Appendix C Linearised massive gravity theory

### v2

@@ Appendix A Effective mass generation of spin-2 field in arbitrary curved spacetime
where $\bar{g}_{\mu\nu}$ is the metric describing the background geometry w.r.t which we linearised the theory of gravity, $h_{\mu\nu}$ is the metric perturbation and $\varepsilon(=2\kappa)$ carries the order of the metric perturbation. This leads to the following expression of the inverse metric:
@@
$\begin{split}\Gamma_{\ \mu\nu}^{\rho}&=\bar{\Gamma}_{\ \mu\nu}^{\rho}+\varepsilon\Big{[}\frac{1}{2}\bar{g}^{\rho\lambda}(\partial_{\mu}h_{\lambda\nu}+\partial_{\nu}h_{\mu\lambda}-\partial_{\lambda}h_{\mu\nu})\\ &-\bar{g}^{\rho\alpha}h_{\alpha\beta}\bar{\Gamma}_{\ \mu\nu}^{\beta}\Big{]}\\ &-\varepsilon^{2}\Big{[}\frac{1}{2}\bar{g}^{\beta\lambda}h_{\ \beta}^{\rho}(\partial_{\mu}h_{\lambda\nu}+\partial_{\nu}h_{\mu\lambda}-\partial_{\lambda}h_{\mu\nu})\\ &-\bar{g}^{\alpha\beta}h_{\ \beta}^{\rho}h_{\alpha\sigma}\bar{\Gamma}_{\ \mu\nu}^{\sigma}\Big{]},\end{split}$ (29)
@@
$\begin{split}*&=(\Gamma_{\ \sigma\nu}^{\beta})_{L}\Big{[}\bar{\nabla}_{\rho}(\bar{g}^{\mu\alpha}h_{\alpha\beta})-\frac{1}{2}\bar{g}^{\mu\alpha}(\bar{\nabla}_{\rho}h_{\alpha\beta}+\bar{\nabla}_{\beta}h_{\rho\alpha}\\ &-\bar{\nabla}_{\alpha}h_{\rho\beta})\Big{]}\\ &=(\Gamma_{\ \sigma\nu}^{\beta})_{L}\Big{[}\frac{1}{2}\bar{g}^{\mu\alpha}(\bar{\nabla}_{\rho}h_{\alpha\beta}+\bar{\nabla}_{\alpha}h_{\rho\beta}-\bar{\nabla}_{\beta}h_{\rho\alpha})\Big{]}\\ &=(\Gamma_{\ \sigma\nu}^{\beta})_{L}\Big{[}\frac{1}{2}\bar{g}^{\mu\alpha}\delta_{\beta}^{\ \lambda}(\bar{\nabla}_{\rho}h_{\alpha\lambda}+\bar{\nabla}_{\alpha}h_{\rho\lambda}-\bar{\nabla}_{\lambda}h_{\rho\alpha})\Big{]}\\ &=(\Gamma_{\ \sigma\nu}^{\beta})_{L}\Big{[}\frac{1}{2}\bar{g}^{\mu\alpha}\bar{g}_{\beta\gamma}\bar{g}^{\gamma\lambda}(\bar{\nabla}_{\rho}h_{\alpha\lambda}+\bar{\nabla}_{\alpha}h_{\rho\lambda}-\bar{\nabla}_{\lambda}h_{\rho\alpha})\Big{]}\\ &=\bar{g}^{\mu\alpha}\bar{g}_{\beta\gamma}(\Gamma_{\ \sigma\nu}^{\beta})_{L}(\Gamma_{\ \rho\alpha}^{\gamma})_{L}.\end{split}$ (47)
@@
$\begin{split}R&=\bar{R}+\varepsilon[\bar{g}^{\nu\sigma}(R_{\nu\sigma})_{L}-h^{\nu\sigma}\bar{R}_{\nu\sigma}]-\varepsilon^{2}\{\bar{g}^{\nu\sigma}h_{\ \beta}^{\mu}(R_{\ \nu\mu\sigma}^{\beta})_{L}\\ &+h^{\nu\sigma}(R_{\nu\sigma})_{L}-h^{\nu\lambda}h_{\ \lambda}^{\sigma}\bar{R}_{\nu\sigma}+\bar{g}^{\nu\sigma}\bar{g}^{\mu\alpha}\bar{g}_{\beta\gamma}[(\Gamma_{\ \sigma\nu}^{\beta})_{L}(\Gamma_{\ \mu\alpha}^{\gamma})_{L}\\ &-(\Gamma_{\ \mu\nu}^{\beta})_{L}(\Gamma_{\ \sigma\alpha}^{\gamma})_{L}]\}\\ &=\bar{R}+\varepsilon\Big{[}\frac{1}{2}\bar{g}^{\nu\sigma}(\bar{\nabla}_{\mu}\bar{\nabla}_{\sigma}h_{\ \nu}^{\mu}+\bar{\nabla}_{\mu}\bar{\nabla}_{\nu}h_{\ \sigma}^{\mu}-\bar{\Box}h_{\sigma\nu}-\bar{\nabla}_{\sigma}\bar{\nabla}_{\nu}h)\\ &-h^{\nu\sigma}\bar{R}_{\nu\sigma}\Big{]}-\varepsilon^{2}\Bigg{[}\frac{1}{2}\bar{g}^{\nu\sigma}h_{\ \mu}^{\rho}[\bar{\nabla}_{\rho}\bar{\nabla}_{\sigma}h_{\ \nu}^{\mu}+\bar{\nabla}_{\rho}\bar{\nabla}_{\nu}h_{\ \sigma}^{\mu}-\bar{\nabla}_{\rho}\bar{\nabla}^{\mu}h_{\sigma\nu}\\ &-\bar{\nabla}_{\sigma}\bar{\nabla}_{\rho}h_{\ \nu}^{\mu}-\bar{\nabla}_{\sigma}\bar{\nabla}_{\nu}h_{\ \rho}^{\mu}+\bar{\nabla}_{\sigma}\bar{\nabla}^{\mu}h_{\rho\nu}]\\ &+\frac{1}{2}h^{\nu\sigma}(\bar{\nabla}_{\mu}\bar{\nabla}_{\sigma}h_{\ \nu}^{\mu}+\bar{\nabla}_{\mu}\bar{\nabla}_{\nu}h_{\ \sigma}^{\mu}-\bar{\Box}h_{\sigma\nu}-\bar{\nabla}_{\sigma}\bar{\nabla}_{\nu}h)\\ &-h^{\nu\lambda}h_{\ \lambda}^{\sigma}\bar{R}_{\nu\sigma}+\frac{1}{4}\Big{[}4\bar{\nabla}_{\sigma}h_{\rho}^{\ \sigma}\bar{\nabla}_{\mu}h^{\rho\mu}+2\bar{\nabla}^{\mu}h^{\lambda\sigma}\bar{\nabla}_{\lambda}h_{\sigma\mu}\\ &-4\bar{\nabla}_{\sigma}h^{\lambda\sigma}\bar{\nabla}_{\lambda}h+\bar{\nabla}_{\rho}h\bar{\nabla}^{\rho}h-3\bar{\nabla}_{\rho}h_{\mu\nu}\bar{\nabla}^{\rho}h^{\mu\nu}\Big{]}\Bigg{]}.\end{split}$ (56)
@@
$\begin{split}\sqrt{-g}R&=\sqrt{-\bar{g}}\Bigg{[}\bar{R}+\varepsilon\left(\frac{1}{2}h\bar{R}-h^{\nu\sigma}\bar{R}_{\nu\sigma}\right)+\varepsilon\underbrace{\Big{[}\bar{\nabla}_{\mu}\bar{\nabla}_{\nu}h^{\mu\nu}-\bar{\Box}h\Big{]}}_{\text{surface terms}}\\ &+\frac{\varepsilon^{2}}{2}h\Big{[}\bar{\nabla}_{\mu}\bar{\nabla}_{\nu}h^{\mu\nu}-\bar{\Box}h\Big{]}\\ &-\frac{\varepsilon^{2}}{4}\left(h_{\alpha\beta}h^{\alpha\beta}-\frac{1}{2}h^{2}\right)\bar{R}+\varepsilon^{2}\bar{R}_{\nu\sigma}\left(h^{\nu\lambda}h_{\ \lambda}^{\sigma}-\frac{1}{2}hh^{\nu\sigma}\right)\\ &-\varepsilon^{2}\left(\frac{1}{4}\bar{\nabla}_{\lambda}h_{\mu\nu}\bar{\nabla}^{\lambda}h^{\mu\nu}+\frac{1}{4}\bar{\nabla}_{\rho}h\bar{\nabla}^{\rho}h-\frac{1}{2}\bar{\nabla}^{\nu}h^{\rho\mu}\bar{\nabla}_{\mu}h_{\rho\nu}\right)\Bigg{]}\\ &=\sqrt{-\bar{g}}\Bigg{[}\bar{R}-\varepsilon h^{\nu\sigma}\bar{G}_{\nu\sigma}-\frac{\varepsilon^{2}}{4}\left(h_{\alpha\beta}h^{\alpha\beta}-\frac{1}{2}h^{2}\right)\bar{R}\\ &+\varepsilon^{2}\bar{R}_{\nu\sigma}\left(h^{\nu\lambda}h_{\ \lambda}^{\sigma}-\frac{1}{2}hh^{\nu\sigma}\right)-\varepsilon^{2}\Big{[}\frac{1}{4}\bar{\nabla}_{\lambda}h_{\mu\nu}\bar{\nabla}^{\lambda}h^{\mu\nu}\\ &-\frac{1}{4}\bar{\nabla}_{\rho}h\bar{\nabla}^{\rho}h-\frac{1}{2}\bar{\nabla}^{\nu}h^{\rho\mu}\bar{\nabla}_{\mu}h_{\rho\nu}+\frac{1}{2}\bar{\nabla}_{\mu}h\bar{\nabla}_{\nu}h^{\mu\nu}\Big{]}\Bigg{]}.\end{split}$ (58)
Now considering $\varepsilon=\kappa$ and rescaling the action of spin-2 field by a factor of 4 (this is equivalent to taking $\varepsilon=2\kappa$ and considering the vacuum Einstein field equations), the linearised Einstein-Hilbert action (without matter) in arbitrary curved spacetime can be expressed as
$\begin{split}S_{EH}&=\frac{1}{2\kappa^{2}}\int d^{4}x\sqrt{-g}R=\frac{1}{2\kappa^{2}}\int d^{4}x\sqrt{-\bar{g}}\bar{R}\\ &+\int d^{4}x\sqrt{-\bar{g}}\Big{[}-\frac{1}{2}\bar{\nabla}_{\lambda}h_{\mu\nu}\bar{\nabla}^{\lambda}h^{\mu\nu}+\frac{1}{2}\bar{\nabla}_{\rho}h\bar{\nabla}^{\rho}h\\ &+\bar{\nabla}^{\nu}h^{\rho\mu}\bar{\nabla}_{\mu}h_{\rho\nu}-\bar{\nabla}_{\mu}h\bar{\nabla}_{\nu}h^{\mu\nu}-\frac{1}{2}\bar{R}\left(h_{\alpha\beta}h^{\alpha\beta}-\frac{1}{2}h^{2}\right)\\ &+2\bar{R}_{\nu\sigma}\left(h^{\nu\lambda}h_{\ \lambda}^{\sigma}-\frac{1}{2}hh^{\nu\sigma}\right)\Big{]}.\end{split}$ (59)
@@
$\begin{split}S_{M}[\Phi,\Phi^{*},g_{\mu\nu}]&=S_{M}[\Phi,\Phi^{*},\bar{g}_{\mu\nu}]+\kappa\int\sqrt{-\bar{g}}h_{\mu\nu}T^{\mu\nu}\\ &-\kappa^{2}\int d^{4}x\sqrt{-\bar{g}}h^{\mu\nu}h_{\rho\sigma}\mathscr{O}_{\mu\nu}^{\ \ \rho\sigma}+\ldots.\end{split}$ (60)
We may note that the linear part cancels with the second term in the first line of the equation (58) since this is nothing but the Einstein’s equation in the presence of matter source. Moreover, it is mathematically consistent since if we expand the total matter and Einstein-Hilbert action around a solution of classical equation of motion, the linear terms drop out and leading order terms are quadratic order in the perturbations.
If we rescale again the spin-2 field contribution by a factor of $4$ coming from the matter sector, we obtain the following total action in the spin-2 field in curved spacetime
$\begin{split}S_{h}&=\int d^{4}x\sqrt{-\bar{g}}\Big{[}-\frac{1}{2}\bar{\nabla}_{\lambda}h_{\mu\nu}\bar{\nabla}^{\lambda}h^{\mu\nu}+\frac{1}{2}\bar{\nabla}_{\rho}h\bar{\nabla}^{\rho}h\\ &+\bar{\nabla}^{\nu}h^{\rho\mu}\bar{\nabla}_{\mu}h_{\rho\nu}-\bar{\nabla}_{\mu}h\bar{\nabla}_{\nu}h^{\mu\nu}-\frac{1}{2}\bar{R}\left(h_{\alpha\beta}h^{\alpha\beta}-\frac{1}{2}h^{2}\right)\\ &+2\bar{R}_{\nu\sigma}\left(h^{\nu\lambda}h_{\ \lambda}^{\sigma}-\frac{1}{2}hh^{\nu\sigma}\right)\Big{]}-\kappa^{2}\int d^{4}x\sqrt{-\bar{g}}h^{\mu\nu}h_{\rho\sigma}\mathscr{O}_{\mu\nu}^{\ \ \rho\sigma},\end{split}$ (61)
@@
where $V(|\Phi_{0}|)$ turns out to be negative after spontaneous symmetry breaking of the field potential. Substituting all the above information including the fact $\bar{R}=-\kappa^{2}T$ and $\bar{R}_{\nu\sigma}=\kappa^{2}\left(T_{\nu\sigma}-\frac{1}{2}\bar{g}_{\nu\sigma}T\right)$ where $T=-4V(|\Phi_{0}|)$ and $T_{\nu\sigma}=-\bar{g}_{\nu\sigma}V(|\Phi_{0}|)$ in the leading order, we obtain the following quadratic action of spin-2 field
$\begin{split}S_{h}&=\int d^{4}x\sqrt{-\bar{g}}\Big{[}-\frac{1}{2}\bar{\nabla}_{\lambda}h_{\mu\nu}\bar{\nabla}^{\lambda}h^{\mu\nu}+\frac{1}{2}\bar{\nabla}_{\rho}h\bar{\nabla}^{\rho}h+\bar{\nabla}^{\nu}h^{\rho\mu}\bar{\nabla}_{\mu}h_{\rho\nu}\\ &-\bar{\nabla}_{\mu}h\bar{\nabla}_{\nu}h^{\mu\nu}-2\kappa^{2}V(|\Phi_{0}|)\left(h_{\alpha\beta}h^{\alpha\beta}-\frac{1}{2}h^{2}\right)\\ &+2\kappa^{2}V(|\Phi_{0}|)\left(h_{\alpha\beta}h^{\alpha\beta}-\frac{1}{2}h^{2}\right)\Big{]}\\ &+\kappa^{2}V(|\Phi_{0}|)\int d^{4}x\sqrt{-\bar{g}}\left(h_{\mu\nu}h^{\mu\nu}-\frac{1}{2}h^{2}\right)\\ &=\int d^{4}x\sqrt{-\bar{g}}\Big{[}-\frac{1}{2}\bar{\nabla}_{\lambda}h_{\mu\nu}\bar{\nabla}^{\lambda}h^{\mu\nu}+\frac{1}{2}\bar{\nabla}_{\rho}h\bar{\nabla}^{\rho}h+\bar{\nabla}^{\nu}h^{\rho\mu}\bar{\nabla}_{\mu}h_{\rho\nu}\\ &-\bar{\nabla}_{\mu}h\bar{\nabla}_{\nu}h^{\mu\nu}+\kappa^{2}V(|\Phi_{0}|)\left(h_{\alpha\beta}h^{\alpha\beta}-\frac{1}{2}h^{2}\right)\Big{]}.\end{split}$ (66)
Appendix B Fierz-Pauli action: Standard, Generalized and Extended